OpenStudy (anonymous):

If m and p are positive integers and (m + p) x m is even, which of the following must be true? A. If m is odd, then p is odd. B. If m is odd, then p is even. C. If m is even, then p is even. D. If m is even, then p is odd.

OpenStudy (amistre64):

distribute, that should help determine the results

OpenStudy (anonymous):

lets see.. (A) m = 3 and p = 5 (m + p) x m = (3 + 5) * 3 = 8 * 3 = 24 (result is even) lets try another m = 7 and p = 9 (m + p) * m = (7 + 9) * 7 = 16 * 7 = 112 (result is even) so what does that tell you

OpenStudy (amistre64):

trial and error, another well established method :)